payton:

is -3(x-4)=-3x+12 a... one solution no solution infinitely many sloutions

darkknight:

what do you think? A first step would be to distribute the left side of the Equation

payton:

distribute -3 by x and -4

darkknight:

yep

darkknight:

what did you get?

ramen:

xs and numbers

payton:

ok so i have -3x+12=-3x+12

darkknight:

-3(x-4)=-3x+12 you distribute to -3x+12 = -3x+12, If you notice these equations are the same on both sides of the equation. So no matter what value you plug in for x you will get a correct statement. Try plugging in like 3 or 7 for x.

darkknight:

Whenever both sides of the equations are the same. Then they have infinite solutions

payton:

okay

payton:

thanks

darkknight:

np, I am the darkknight